OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x^2 - 2x y = -2x^2 + 4x - 3

hero (hero):

y = x^2 - 2x opens upward y = -2x^2 + 4x - 3 opens downward use x = -b/2a to find the x coordinate of vertex, then insert that into original parabola to find y